OpenStudy (anonymous):

5x2+7x+2 factor as a polynomial in descending order help!!!!!!!!!!!!!!

OpenStudy (mathstudent55):

There is no common factor we can factor out. Here we have again a trinomial in which the first term is not just x^2, but 5x^2. We use the ac method. ac = 5 * 2 = 10 Now we need two factors of 10 that add to 7.

OpenStudy (anonymous):

5,2

OpenStudy (mathstudent55):

Good. Now we break up the middle term: \(5x^2 + 2x + 5x + 2 \) Now we factor by parts: \(x(5x + 2) + 1(5x + 2) \)

OpenStudy (mathstudent55):

Now we factor out the commn factor: \( (5x + 2)(x + 1) \)

OpenStudy (anonymous):

what about the factorization of -2x3+2x2+12x

OpenStudy (mathstudent55):

Pls start a new post for each problem.